OpenStudy (anonymous):

The ratio of the width to the length for the rectangle is JK:KL is 2:3 Solve for the variable of the length: z-3, given width( JK ) = 12

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 } = \frac{ 12 }{ z - 3 }\] \[z-3 = \frac{ 12*3 }{ 2 }\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

@smartkid2000 whats your question?